40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 26 17:50:22 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
    Reply  Wed Jan 27 18:19:45 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
       Reply  Thu Jan 28 00:29:30 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement PLL_OLG.pdfFreq_noise_comparison.pdf
          Reply  Mon Feb 1 18:11:25 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement Free_running_frequency_noise_comparison.pdf
       Reply  Fri Jan 29 14:43:47 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.pdf
          Reply  Fri Jan 29 16:55:27 2016, rana, Update, Green Locking, Innolight laser is 10 years old 
             Reply  Sat Jan 30 00:12:22 2016, gautam, Update, Green Locking, Innolight laser is 10 years old RIN_comparison.pdf
                Reply  Sat Jan 30 09:56:24 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                   Reply  Mon Feb 1 15:16:28 2016, gautam, Update, Green Locking, Innolight laser is 10 years old 
                      Reply  Mon Feb 1 15:43:18 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                         Reply  Fri Feb 5 15:02:13 2016, gautam, Update, Green Locking, X-end NE cable 
          Reply  Fri Feb 12 11:10:49 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.jpg
          Reply  Mon Mar 21 14:29:32 2016, Steve, Update, Calibration-Repair, 1W Innolight laser repair diagnoses 
             Reply  Thu Mar 24 07:56:09 2016, Steve, Update, Calibration-Repair, NO Noise Eater for 1W Innolight 
                Reply  Mon Apr 11 17:03:41 2016, Steve, Update, Calibration-Repair, 1W Innolight repair completed 
Message ID: 12045     Entry time: Thu Mar 24 07:56:09 2016     In reply to: 12040     Reply to this: 12070
Author: Steve 
Type: Update 
Category: Calibration-Repair 
Subject: NO Noise Eater for 1W Innolight 

1W Innolight is NOT getting Noise Eater as it was decided yesterday at the 40m meeting. Corrected 3-25-2016

Repair quote with adding noise eater is in 40m wiki

Quote:

 

Quote:
Quote:

After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...I

Innolight 1W 1064nm, sn 1634 was purchased in 9-18-2006 at CIT. It came to the 40m around 2010

It's diodes should be replaced, based on it's age and performance.

RIN and noise eater bad. I will get a quote on this job.

The Innolight Manual frequency noise plot is the same as Lightwave' elog 11956

Diagnoses from Glasglow:

“So far we have analyzed the laser. The pump diode is degraded. Next we would replace it with a new diode. We would realign the diode output beam into the laser crystal. We check all the relevant laser parameters over the whole tuning range. Parameters include single direction operation of the ring resonator, single frequency operation, beam profile and others. If one of them is out of spec, then we would take actions accordingly. We would also monitor the output power stability over one night. Then we repackage and ship the laser.”

 

ELOG V3.1.3-